OpenStudy (anonymous):

On a Wisconsin road map, Kristin measured a distance of 11 1/8 (11.125) in. from Madison to Wausau. The scale of this map is 1 inch for every 13 miles. What is the actual distance between Madison and Wausau to the nearest mile?

OpenStudy (johnweldon1993):

okay well lets see if 11.125 inches...is the distance between the 2....and you want that in miles and you know that scale is 1 inch = 13 miles \[\frac{ 1 }{ 13 } = \frac{ 11.125 }{ x }\] so you would multiply both sides by x.....and them multiply both sides by 13 which will leave you with 1x = 11.125(13) so 11.125 times 13...is the distance in miles between the 2....which is?

OpenStudy (anonymous):

144.6 so 145mi? right (: ?

OpenStudy (johnweldon1993):

that would be correct :)

OpenStudy (anonymous):

thanks ♥

OpenStudy (johnweldon1993):

anytime :)